trying to help my daughter work this problem in pre-algebra 6m-2 1/2 = m+ 12 1/2. I have tried several different ways, but have been unsucessful
Answer by Alan3354(69443)

6m-2 1/2 = m+ 12 1/2
Multiply thru by 2 to eliminate fractions
12m - 5 = 2m + 25
Add 5
12m = 2m + 30
Subtract 2m
10m = 30
m = 3